OpenStudy (anonymous):

How many multiples of 4 are there between 12 and 96, inclusive. Show work please? Found this in my SAT study guide

OpenStudy (anonymous):

I got 21 but not sure if it is correct

OpenStudy (anonymous):

96/4 = 24 12/4 = 3 4 and 8 are excluded 24 - 2 = 22 it should be 22

OpenStudy (anonymous):

Aiming for MIT or Harvard?

OpenStudy (anonymous):

It is my dream , a very far away dream :P

OpenStudy (anonymous):

good luck :)

OpenStudy (anonymous):

What about yourself? I started caring about grades pretty late, so I'm sure it is too late :(

OpenStudy (anonymous):

I'm still in Grade 10. I'm aiming for one of the universities in Canada

OpenStudy (anonymous):

Oh awesome! Just take advantage of everything you have. There's so much free knowledge on the internet but most people use it for social networking. Good luck with your dreams. Hope you achieve them

OpenStudy (anonymous):

it's not really my dream, but it's my plan

OpenStudy (anonymous):

Oh nice ;)